OpenStudy (anonymous):

Please help: x+1/x+2=x+3/x-1=1

OpenStudy (anonymous):

solve for x

OpenStudy (shadowlegendx):

\[x + 1 \div x + 2 = x + 3 \div x - 1 = 1\]

OpenStudy (shadowlegendx):

@januaryjoy29 what is the first step you do when solving for a variable?

OpenStudy (shadowlegendx):

If you don't know, that's completely fine :)

OpenStudy (anonymous):

sorry i dont know, clueless

OpenStudy (shadowlegendx):

Ok don't worry!

OpenStudy (shadowlegendx):

1st step you would do is isolate the variable

OpenStudy (shadowlegendx):

This basically means to get it alone so you can have something like: x = 1 or x = 7 Those are not the answers, just examples :)

OpenStudy (anonymous):

im trying to understand

OpenStudy (shadowlegendx):

Don't worry I'm gonna try explain this the best I can because I remember when I had trouble with Algebra :P

OpenStudy (anonymous):

ok thanks

OpenStudy (shadowlegendx):

Sorry I got disconnected from the site, it does that sometimes :D

OpenStudy (anonymous):

The problem involves: Solving a rational equation that simplifies to quadratic: Binomial denominators and numerators

OpenStudy (shadowlegendx):

\[x + 1 \div x + 2 = x + 3 \div x - 1 = 1\] Add 1 to the negative 1 and the positive two You get \[x + 1 \div x + 3 = x + 3 \div x = 1\]

OpenStudy (anonymous):

the answer is suppose to be two numbers like:0,6

OpenStudy (shadowlegendx):

\[x+1÷x+3=x+3÷x=1 \] Subtract 3 from the positive 3 on both sides You get \[x + 1 \div x = x \div x = 1\]

OpenStudy (anonymous):

so your saying the answer is 1?

OpenStudy (anonymous):

ok thanks s much

OpenStudy (shadowlegendx):

The answer is not 1, we are solving for x

OpenStudy (shadowlegendx):

Can you post you answer choices?

OpenStudy (shadowlegendx):

@januaryjoy29 Can you post your answer choices?

OpenStudy (shadowlegendx):

Ok so when we come to the point \[x \div x = x \div x = 1\] We see that this equation is true. So x does equal 1. The reason why I said x wasn't 1 earlier is because solving for equations early when you think you know the answer CAN be disastrous and I discourage it. Though when it is clear as this, you need not solve further unless your teacher tells you to do so.

OpenStudy (anonymous):

its the Aleks system yo have to figure it out

OpenStudy (shadowlegendx):

Aleks system?

OpenStudy (anonymous):

yes

OpenStudy (shadowlegendx):

Haha what is that?

OpenStudy (anonymous):

online system..thank you for all your help